Recent Log Entries

Autolog save

Log Number: 3829868Logbook(s): Tags: Body: saver ran from opsl632

Saved: ywang 08-18-20 10:03:39 {A:1 B:5 C:5 E:1050 A=Tune B=BS C=BS D=BS 0R08=0.00 0R08 dump in. A B C dump(s) in.

Injector optimized

Log Number: 3829867Logbook(s): Entry Makers: ywangBody: 

0850 - 1000:
Fig. 1 and Fig. 2 show 130uA and 150uA to ILD and Fcup #2, respectively, after optimization.
Waitint for Hall A to check PQB.

IHAGT06A - preamp check out

Log Number: 3829862Logbook(s): Entry Makers: delacruzBody: 

Singal injected the preamp with a 1mV @ 100mHz signal, measured the output directly from the preamp and verified the voltage output. Verified ADC signal read back on the expert control screen (see fig. 2 and 3).

Cycled harp, recorded the harp scan (see fig.1).

*During the troubleshooting process, I was measuring -5.0 VDC on the preamp output. After disconnecting the signal cable from the harp to the preamp and then reconnecting them, the negative voltage disappeared.

Checking on Hall B Beam Noise in Injector

Log Number: 3829845Logbook(s): Entry Makers: suleimanBody: 

Very quickly tried to look for noise on Hall B beam in the Injctor Service Building. Nominal B beam was on to FC2 and looking at BPM 0I02. Figure one shows the setup. It seems there is motion on the beam at 11 kHz (Figure 2). Zooming in, Figures 3 and 4 show two peaks at 10.46 and 10.78 kHz. No such noise on 1I02 - Figure 5.

Once Yan is done steering the injector, will check again for noise on Hall B beam.

COMMENT ON MBSY3C inspection of PSS interlock signal connector at box supply

Log Number: 3829840Logbook(s): Body: 

Task discussed with owner and approved.
Inspection and/or repair of the connector will require functional tests of the interface. Paul Metcalf will assist with certification. Crew Chief will need to direct the certs as usual.
Certification will take around 15 min. and should be performed when BSY and Hall C can be configured up to Beam Permit PSS State (most likely towards the end of the maintenance period).

IPM2L24 has been disabled

Log Number: 3829836Logbook(s): Body: Problem: Calibration Failure
BPM broken during CW beam. Ran calibration and it fails Calibration. I think the RF module got irradiated. Disabling this bpm and its associated passes.

Pages